Solve: 1 / 18 + 1 / 18 Solve: 1 / 18 + 1 / 18 Solve: 1 / 18 + 1 / 18

Solve:
1 / 18 + 1 / 18
Solve:
1 / 18 + 1 / 18
Solve:
1 / 18 + 1 / 18

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

1 / 18 + 1 / 18
= (1 + 1) / 18
= 2 / 18
=1 / 9

​Thanks
Shraddha Singh
2 years ago
Dear Student,
 
The denominator of both the fraction is same, so , we need not to take LCM.
1/18 + 1/18
= (1+1)/18
= 2/18
= 1/9
 
Hope it helps!
Thank you